### v1

@@ 4.2.2 Performance comparison with other reward functions
In this experiment, we conducted trials applying the reward functions from studies [13, 16, 31] introduced in Section II-B, along with the reward function incorporating TGRF described in Section III-B-3. The navigation method employed GST + HH Attn models. Participants adhered to the ORCA approach, recording SR, NT, PL, ITR every 2000 episodes. The total episodes conducted were 20000 for GST + HH Attn.
First, it was observed that applying TGRF to the model led to an overall performance improvement compared to other models. In Fig. 8a, the reward function with TGRF was able to drive the algorithm’s performance up to a maximum of 95% over 16000 total episodes. Conversely, other reward functions achieved a maximum of 90% SR. This indicates that TGRF harmonizes appropriately with other reward functions, assuming a shape that aligns with the role of rewards, thereby eliciting the algorithm’s maximum performance.
In Fig. 8b, it was noticed that in the model applying TGRF, NT decreased with repeated learning, ultimately confirming the second lowest NT. Correspondingly, in Fig. 8c, the second lowest PL was also observed. This is associated with ITR, as lower NT and PL imply the robot tolerating penalties due to $r_{disc}(s_{t})$ to reach the destination, resulting in higher ITR. For instance, looking at the orange graph, the highest ITR along with the lowest NT and PL can be observed. As a result, in Fig. 7a, this leads to choosing shorter and riskier paths, increasing the likelihood of not understanding human intentions and a higher possibility of collisions. However, in Fig. 8d, the model incorporating TGRF maintains the lowest ITR in most cases. This demonstrates that by applying TGRF, it selects the most efficient and safe paths compared to other models while maintaining the highest SR, reflecting the intentions of the algorithm, as shown in the results of Fig. 7b.
Second, Fig. 8 demonstrated the significant advantage of TGRF in terms of learning speed compared to the reward function without TGRF. In Fig. 8a, it was observed that three models reached saturation from a total of 6000 episodes. At this point, the model applying TGRF achieved the highest SR. This indicates the contribution of TGRF to faster learning speeds.
However, the TGRF has limitation in crowded environments. It does not inherently enhance the performance of core algorithms. For certain policies, comparable performances were achieved, as shown in Tables I and II. This suggests that the TGRF expedites the algorithm to achieve an optimal performance rather than inherently enhancing the algorithm itself.
@@ 5 Real-World Experiments
Figure 9: Evasive actions performed by the robot in real-world scenarios and corresponding renderings with four humans (from left to right). The green arrows represent the movement path of the human, while the red arrows indicate the movement path of the robot. In addition, yellow circles indicate robots, green circles represent humans, and red stars indicate the destinations. These illustrations showcase the avoidance strategies employed by the robot as it encounters successive individuals: the (a) first, (b) second, (c) third, and (d) fourth human.
@@
This study extends beyond simulations to real-world experiments. The model, trained using the unicycle approach, was applied to a physical robot in a real environment. Unicycle kinematics present limitations in terms of direction and position control as compared to holonomic kinematics. Our experimental setup comprised a host computer equipped with an Intel i5-10600 processor @ 3.30 GHz and Nvidia RTX 3070 GPU; the robot utilized was a Turtlebot3. A LiDAR sensor, LDS-01, played a pivotal role in the human detection and robot position estimation. Robot positioning relied on SLAM localization, and human detection was accomplished using a 2D people detection algorithm based on 2D LiDAR data [35]. Detailed information can be accessed from https://github.com/JinnnK/Tranformable-Gaussian-Reward-sim2real-with-Turtlebot3.
@@
Although our study assumed an absence of static obstacles other than humans, our real-world experiments were conducted in a confined space measuring approximately 3 × 5m with static obstacles. These experiments involved scenarios in which the robot navigated between predefined start and destination points, encountering one to four pedestrians along its path. The maximum speed was approximately 0.6 m/s, and the investigation covered scenarios involving four moving individuals.
@@
In Fig. 9a, the robot facing diagonally upward as the pedestrian moves from right to left. In this scenario, the robot rotates to the left, aligning with the pedestrian’s movement direction, instead of moving behind (to the right of) the human. This decision appears rational because both the destination and robot’s current orientation are oriented diagonally upward, making a leftward maneuver the most efficient choice when considering the human direction, speed, and destination. Notably, in another experiment involving three individuals, the robot was observed to halt temporarily instead of moving to the left.
@@
In Fig. 9b, the robot encounters a pedestrian walking diagonally from left to right. In response, the robot navigates to the left to avoid obstructing the path of the pedestrian.
@@
In Fig. 9c, the robot faces a human walking from left to right. Similarly, it predicts the human trajectory and executes a leftward turn to avoid a collision while approaching the destination.
@@
In Fig. 9d, the robot encountered a human crossing diagonally from left to right near the destination. The robot smartly avoided the human by initially turning to the left, avoiding the pedestrian, and then turning to the right to reach the destination.
@@
These actions underscore the robot’s ability to make real time decisions based on a dynamic environment, considering factors such as the human path, velocity, and proximity to the destination. The robot’s avoidance strategies prioritize efficiency while maintaining safety, and are influenced by various factors, including its current orientation and overall context of the situation. These real-world experiments testify the model’s adaptability in complex and dynamic environments, where human-robot interactions demand responsive and context aware behavior. The comprehensive renderings and additional experimental videos are available on https://youtu.be/9x24k75Zj5k?si=OtczdVXPUnbGwpv-.
@@
Two primary limitations were encountered during this experiment. (1) Computation load: the use of DNNs for action and trajectory predictions significantly increases computational demands. Considering the number of pedestrians considered, particularly for trajectory prediction, the time required for the next action was approximately 0.22 s. This resulted in seemingly irregular robot movements and delayed pedestrian responses. (2) Physical constraints: the accuracy of human detection and prediction is affected by sensor noise, limitations in the detection performance, and challenges in determining human angles. These factors lead to the occasional misidentification of obstacles as humans or limitations in the precision of human location information, thereby reducing the accuracy of the trajectory prediction. In addition, noise from the LiDAR sensor and location information errors caused by the robot’s movement accumulated over time, resulting in inaccuracies in location values as the experiment progressed.
6 Conclusion and Further Works
@@
This study proposed the TGRF tailored for robots navigating crowded environments. This reward function has three advantages: a good performance with few hyperparameters, versatility for various objectives, and faster learning and stabilization. These statement were improved by the success rates and algorithm’s ability to discern human intentions compared with previous reward functions.
@@
However, challenges emerged in both the simulations and real-world experiments. In the simulations, these challenges involved sensitivity to hyperparameters, algorithmic limitations, a trade-off correlation between the SR and NT, and absence of static obstacles. In real-world tests, challenges include the sensor noise, and physical constraints.
@@
Hence, in future research, (1) we will apply TGRF in various environments and with different objects. In this study, we demonstrated its performance limited to rewards concerning humans. Therefore, we plan to conduct experiments by applying TGRF to diverse objects such as walls, obstacles, drones, and more. (2) we intend to devise TGRF considering physical limitations. While TGRF performs exceptionally well in ideal conditions, its performance decreases in reality due to computational load and physical issues. Hence, we aim to implement a dynamically adaptive TGRF that adjusts based on the situation by incorporating knowledge regarding these physical limitations.

### v2

@@ 4.2.2 Performance Comparison with Other Reward Functions
In this experiment, we conducted trials applying the reward functions from previous studies [13, 16, 31] introduced in Section II-B along with the reward function incorporating the TGRF described in Section III-B-3. The navigation method employs the GST + HH Attn models. Participants adhered to the ORCA approach by recording SR, NT, PL, and ITR every 2000 episodes. The total episodes conducted were 20000 for GST + HH Attn.
First, it was observed that applying TGRF to the model led to an overall performance improvement compared with the other models. As shown in Fig. 8a, the reward function with the TGRF was able to drive the algorithm’s performance up to a maximum of 95% over 16000 total episodes. Conversely, the other reward functions achieved a maximum SR of 90% This indicates that the TGRF harmonizes appropriately with the other reward functions, assuming a shape that aligns with the role of the rewards, thereby eliciting the algorithm’s maximum performance.
As depicted in Fig. 8b, in the model applying the TGRF, the NT decreased with repeated learning, ultimately confirm-ing the second lowest NT. Correspondingly, in Fig. 8c, the second lowest PL was also observed. This is associated with the ITR, as lower NT and PL imply that the robot tolerates penalties owing to $r_{disc}(s_{t})$ reaching the destination, resulting in a higher ITR. For instance, in the orange graph, the highest ITR, along with the low-est NT and PL values, can be observed. As a result, as shown in Fig. 7a, this leads to the choice of shorter and riskier paths, increasing the likelihood of not understanding human intentions and a higher possibility of collisions. However, as shown in Fig. 8d, the model incorporating the TGRF maintained the lowest ITR in most cases. This demonstrates that the TGRF selects the most efficient and safe paths compared to the other models, while maintaining the highest SR, reflecting the intentions of the algorithm, as shown in the results of Fig. 7b.
Second, Fig. 8 demonstrates the significant advantage of the TGRF in terms of learning speed compared to the reward function without the TGRF. As shown in Fig. 8a, the three models reached saturation after 6000 episodes. At this point, the model applying TGRF achieved the highest SR. This indicates that the TGRF contributes to faster learning speeds.
However, the TGRF has limitations in crowded environments. It does not inherently enhance the performance of the core algorithms. Comparable performance was achieved for certain policies, as shown in Tables I and II. This suggests that the TGRF expedites the algorithm to achieve optimal performance, rather than enhancing the algorithm itself.
@@ 5 Real-World Experiments
Figure 9: Evasive actions performed by robot in real-world scenarios and corresponding renderings with four humans (from left to right). Green arrows represent the movement path of human, while red arrows indicate the movement path of robot. In addition, yellow circles indicate robots, green circles represent humans, and red stars indicate destinations. These illustrations showcase avoidance strategies employed by the robot as it encounters successive individuals: (a) first, (b) second, (c) third, and (d) fourth human.
@@
This study extended beyond simulations to real-world experiments. The model trained using the unicycle approach was applied to a physical robot in a real environment. Unicycle kinematics presents limitations in terms of direction and position control when compared to holonomic kinematics. Our experimental setup consisted of a host computer equipped with an Intel i5-10600 processor at 3.30 GHz and an NVIDIA RTX 3070 GPU, which was integrated with a Turtle-bot3. A LiDAR sensor, LDS-01, played a pivotal role in human detection and robot position estimation. Robot positioning relies on SLAM localization, and human detection is accomplished using a 2D people detection algorithm based on 2D LiDAR data [35].
@@
Although our study assumed the absence of static obstacles other than humans, our real-world experiments were conducted in a confined space measuring approximately 3 × 5m with static obstacles. These experiments involved scenarios in which the robot navigated between predefined start and destination points and encountered one to four pedestrians along its path. The maximum speed was approximately 0.6 m/s, and the investigation covered scenarios involving four moving individuals.
@@
As shown in Fig. 9a, the robot faces diagonally upward as the pedestrian moves from right to left. In this scenario, the robot rotated to the left, aligning with the pedestrian’s direction of movement, instead of moving behind (to the right) the human. This decision appears rational because both the destination and the robot’s current orientation are oriented diagonally upward, making a leftward maneuver the most efficient choice when considering the human direction, speed, and destination. Notably, in another experiment involving three individuals, the robot was observed halting temporarily instead of moving to the left.
@@
In Fig. 9b, the robot encounters a pedestrian walking diagonally from left to right. In response, the robot navigated to the left to avoid obstructing the path of the pedestrian.
@@
In Fig. 9c, the robot faces a human walking from left to right. Similarly, it predicts the human’s trajectory and executes a leftward turn to avoid collisions while approaching the destination.
@@
In Fig. 9d, the robot encounters a human crossing diagonally from left to right near the destination. The robot smartly avoided humans by initially turning left, avoiding the pedestrian, and then turning right to reach its destination.
@@
These actions underscore the robot’s ability to make real-time decisions based on a dynamic environment, considering factors such as human path, velocity, and proximity to the destination. The robot’s avoidance strategies prioritize efficiency while maintaining safety and are influenced by various factors, including its current orientation and the overall context of the situation. These real-world experiments verified the adaptability of the model in complex and dynamic environments, where human–robot interactions demand responsive and context-aware behavior. Comprehensive renderings and additional experimental videos are available on https://youtu.be/9x24k75Zj5k?si=OtczdVXPUnbGwpv-.
@@
Two primary limitations were encountered during this experiment. First, the computation load: The use of DNNs for action and trajectory predictions significantly increases the computational demands. Considering the number of pedestrians, particularly for trajectory prediction, the time required for the next action was approximately 0.22 s. This resulted in irregular robot movements and delayed pedestrian responses. Second, physical constraints: The accuracy of human detection and prediction is affected by sensor noise, limitations in detection performance, and challenges in determining human angles. These factors lead to occasional misidentification of obstacles as humans or limitations in the precision of human location information, thereby reducing the accuracy of trajectory prediction. In addition, noise from the LiDAR sensor and location information errors caused by the movement of the robot accumulated over time, resulting in inaccuracies in the location values as the experiment progressed.
6 Results and Future Research
@@
This paper introduces a TGRF specifically designed for robots navigating crowded environments. The TGRF offers several advantages, including high performance with minimal hyperparameters, adaptability to diverse objectives, and expedited learning and stabilization processes. These claims are supported by the success rates achieved and the algorithm’s enhanced ability to discern human intentions when compared to previous reward functions.
@@
However, challenges have emerged in both the simulations and real-world experiments. In the simulations, these challenges involved sensitivity to hyperparameters, algorithmic limitations, a trade-off correlation between SR and NT, and the absence of static obstacles. In real-world tests, the challenges include sensor noise and physical constraints.
@@
Hence, in future research, we propose two key strategies, (1) we will apply the TGRF to various environments and with different objects. While our study demonstrates its effectiveness primarily with human rewards, we plan to expand our experiments by applying the TGRF to diverse objects, such as walls, obstacles, and drones. (2) we will devise a TGRF that considers physical limitations. Although the TGRF performs exceptionally well under ideal conditions, its perfor-mance decreases in reality due to computational load and physical issues. Therefore, we aim to implement a dynamically adaptive TGRF that adjusts according to the situation by incorporating knowledge regarding these physical limitations.